INAF LBC+MODS queue observing, 2012 February 18-19

Observer: K. Boutsia, M. Fumana
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: D. Thompson
Instrument: LBC+MODS

Summary

Clear skies at sunset . There was a big chunk of snow on the roof still but after consulting with JohnH we decide to open.

We tried to run the Non-Sidereal progam, but BP14 does not support offsetting within the OB (dither patterns). We have successfully managed to send "stare" OBs in NS mode. The PI has been informed and they should adapt their OBs, before we try to run this program again.

The mirrors were not thermalized, since the telescope had been closed for the past week and it has been really hard to collimate. It took us 1hr at the beginning of the night. For the whole night it would dirft out of focus even immediately after the dofpia. At some point, after doing dofpia, we run the focus OB for rangebal and the red image was already out of collimation. It took a minimum of 5 iterations for dofpia to collimate every time!.
At 5:30UT, the TO realised that the ventilation on the blue side was off. This was contributing to our collimation problems for the first half of the night. Even after though, we did not manage to properly maintain focus for the whole night.
Probably trying to use LBC after the dome has been closed for a week, it is not such a good idea!!
After correcting the ventilation problem, the blue side became much more stable, but the red was constanty drifting out of focus. Most of the times we had to abort the OBs and run dofpia again.

Most of the data we obtained today are out of specs. We have just been going from target to target, trying to collimate.

At 8:51UT we close because of hight humidity.

Data Summary

Project Objects Filter ExpTime Notes
Valsecchi 1999AO10 g-sloan / r-sloan - tried to do Non-siderea tracking but it did not work - we sent suggestions to the PI how to adapt the OBs
Storm IC10VI V-bess / I-bess 10min aborted twice, first because the blue side was out of focus, then because seeing ~1.8'' - fwhmb=2.12 / fwhmr=1.75 - no useful data
Bedin M35_N2158 V-bess / i-sloan 20min first images within spec but after the 2nd image in an OB it could not hold the focus and would go out of spec - fwhmb=1.44 / fwhmr=1.67 - no useful data
Donati be27 B / V,I 12min fwhmB=1.83 / fwhmV=1.59 / fwhmI=1.96 - all data out of specs
Donati be30 B / V,I 12min fwhmB=1.41 / fwhmV=1.54 / fwhmI=1.04 - marginally within specs as far as it concerns seeing and red images elongated: average ellipR=0.13 with peaks of 0.2
Donati ngc2158 B / V,I 12min fwhmB=1.12 / fwhmV=1.37 / fwhmI=1.05 - average ellipticity on the red: 0.11 - this set could be considered within specs.
Gilli field1 rsloan / i-sloan,z-sloan 36min fwhmblue=1.47 / fwhmred=1.54 (there are 12min with fwhmred=1.1") - out of specs

Details

00:30 10Bias_Bino - twice

01:16 Slewing for flats
  • SkyFlat_Vr_pa0 - 0.5 scale - counts 45000-25000Blue / 35000-15000Red
  • SkyFlat_Vr_pa180 - 1.0 scale - counts 26000-15000Blue / 20000-10000Red
  • SkyFlat_gi_pa180 - 3 scale - counts 50000-25000Blue / 26000-15000Red
  • SkyFlat_gi_pa0 - 10 scale - counts 26000-15000Blue / 18000-9000Red - stars are visible

01:51 Use the focusOB 20120219F for dofpia and lbcrangebal
  • dofpia,/First,/X2
-13944      0      0      0      0      0      0   1.40   5.49     red
 -6857      0      0      0      0      0      0   1.80   2.70    blue

-1889   -217    332  -1184    587    508     64   1.54   1.01     red
 2716   -432   -296   -588   -332      0      0   2.54   1.13    blue

 1071   -435    358    -60    475    457     74   2.17   0.58     red
-3325   -314   -398     31   -409    518     64   2.13   1.37    blue

-2019    -26     90    468   -232    261     14   1.47   0.85     red
 2182     47   -178     95   -144     67     -7   2.35   0.87    blue

 1278      9     87   -316    -44    -15     47   1.87   0.53     red
-1174   -111   -219     19   -160    325     99   2.15   0.52    blue

 -2207    161     57    -28    233     -4     -2   1.39   0.88     red
-1526    151   -286      4     74    174     -3   1.77   0.62    blue

-1509      1   -159    341   -146      6     12   1.11   0.62     red
-1009    -11    105   -104     18    128     25   1.64   0.42    blue

 1325   -148    -23   -295     99     74     22   1.65   0.55     red
No GOOD pupils found. - blue

-1281     86   -115    234     28   -107    -13   1.01   0.53     red
 -927     42    -88   -144     98     88     20   1.75   0.39    blue

 1331     79   -138    -97    254      9      2   1.02   0.55     red
 663     98    172    266    -96     62      2   1.81   0.31    blue

 -84     45    -52     60   -205     68     15   1.25   0.14     red
 1201   -132     46    380    355    212     52   2.01   0.55    blue

 -289    -19     68     67    145    -71    -14   1.02   0.16     red
 -1967   -169     15   -275    306     -3      0   1.69   0.80    blue

clear active optics on blue side...
 -1923    -21   -152    912   1718    191     24   2.26   1.16    blue
  • something happened and it does not measure red anymore...
  • temperature difference between mirrors and ambient ~3.5deg!!!
  • there is big gradient also on the mirrors themselves.

02:20 something went wrong when we cleared optics on blue
  • exit previous dofpia after 16th iteration!! - manually put -0.5mm globalfocus on both sides and restart
  • dofpia,/First,/X2
 -6113      0      0      0      0      0      0   0.78   2.41     red
 1020   -289   -366    109    744      0      0   2.73   0.55    blue

  106   -204      1    -33   1035    247      0   1.04   0.50     red
 No GOOD pupils found. - blue

 -1823    -33    110   -878    537    530     41   0.86   0.92     red
 -435   -134   -622     69    232    315     60   1.97   0.35    blue

 82     21    145    589    -79    119     28   0.91   0.29     red
-2243     67   -724   -322  -2715   1000     57   0.00   1.65    blue

  -388    -20    143     80   -101    197     25   0.80   0.22     red
 416   -206     68    -15    566      3    -19   1.72   0.32    blue

 315    -37     59   -209    131     76     18   0.88   0.19     red
 -828    -21    156    228    128    -58      5   1.39   0.36    blue

 353     84     85    267     54    -35    -11   0.69   0.21     red
  61   -154     11     16    -12    -61    -20   1.52   0.10    blue

 334    -17     54    -61    -18    -52     -9   0.77   0.16     red
155    135    247    -72     54    -59    -12   1.37   0.14    blue
  • lbcrangebal to copoint and then dofpia again because we are not convinced by collimation
  • dofpia, /X2
1455     53     84    191      6    -89    -16   0.80   0.59     red
 496    -88    123    225   -170   -142    -18   1.39   0.27    blue

 569     -3    145   -208    241     60     15   1.27   0.28     red
-200     68    120     43     41     41     19   1.79   0.12    blue

 -164    206    -34    467   -151    -22      0   0.94   0.25     red
 52   -116     48    -54    -44    176     12   1.31   0.15    blue

02:41 NSIGUI start - send 20120218E
  • presest cancelled because offsets not allowed...
  • send the preset again with NO_GUIDING!
  • there was on original offset at the preset - to change this we added the original RA,DEC offsets to the dither pattern

02:52 dofpia,/X2
 -1946    -29   -118     99    -33   -111    -24   0.58   0.78     red
  499    -10   -199    121    222   -277    -53   1.29   0.30    blue

 -494      8   -148     64    111    -45     -2   0.67   0.22     red
-346   -116   -180     99    106    -63     -5   1.39   0.19    blue

  -565   -114    -29    -65    158    -15     -6   0.71   0.25     red
 -196   -124   -163     95   -154    -79     -6   1.50   0.16    blue

 -111      6    -91    159    -16     48     -3   0.72   0.12     red
  902   -125    -22    -21     36     38      0   1.54   0.37    blue

02:59 20120218E_again - preset cancelled again!!!
  • offset is not accepted

03:04 set NSIGUI again and send the focusOB to see if it works ... since this one has no offsets!!
  • the preset gave error again and failed - get trajectory failed...
  • it seems that we cannot do this!
  • we give up!

03:10 we have problems with the presets - IRTC seems to be authorized!! - we cannot send presets
  • there is a track preset hanging in the iif gui - call CBiddik

03:22 we slew to the IC10VI field because we have 1.5seeing on the blue and 1.2seeing on the red so Bedin is in principle out of specs...
  • we have no focus OBs and we just slew there and dofpia...
  4698    -93    214   -540    145     85     -7   1.16   1.87     red
   54    178     74    288    138    100    -71   1.18   0.19    blue

  1   -103    -50    307     -8     74      0   0.98   0.17     red
 -17    166    121    162    282      4    -73   1.08   0.18    blue

03:28 IC10VI
  • red images are ok and fwhm is 1.2"
  • blue images are elongated and fwhm is >1.5"
  • we stop... because blue side is horrid!

03:36 dofpia, /X2
-1074    -97    -61     80    187     63     29   1.05   0.44     red
 1265    195    -11    156    203     28    -46   1.13   0.52    blue

 -365    -18   -144     69   -148     52     29   1.15   0.19     red
 169    -51    152    -18     46     90     -5   1.23   0.13    blue

 -1445    -58    143    161    -63    276     44   1.23   0.61     red
  375    119     63    102     80    -39    -24   1.31   0.18    blue

-1047   -103    143    110    -67     72     29   1.20   0.43     red
  85    -79    117    100     51     86     -1   1.36   0.12    blue

  -200    -14    110     32     42    -21      9   1.10   0.12     red
  461    247    111     96    -60     -9    -34   1.16   0.22    blue

03:45 IC10VI
  • seeing on the dimm 2.1'' - measured on the images 1.8-2.0''
  • it does not seem to be an elongation problem anymore...
  • temperature went down by 3deg in the last 2hr.

04:00 seeing measured on images >2.0'' - abort IC10VI
  • go on with NSIGUI testing....

04:01 send the 20120218E_again with no-guiding but again it does not work!!
  • we tried the same procedure with the focus OB and it seems to be woring in stare mode...
  • the PI should send 10OBs - 1 for every offset they want to do.
  • we try offsets in the presets and no-dither in the OB and it seems to be working... 20120218E_nooffset
  • we try once more with guiding on! it also works like this...

04:25 Done with NS tests... slew to Bedin and try our luck!
  • focus OB for Bedin
  • dofpia,/X2
 12510      0      0      0      0      0      0   0.69   4.93     red
 3393   -343    124    472     97   -261     -3   1.00   1.37    blue

04:34 dofpia aborted

04:37 dofpia,/X2

04:38 Aborted again. Focus offset manually changed

04:39 dofpia,/X2
-1449    129     -5  -1239    443    -15   -100   0.96   0.83     red
-953    -75   -104   -139    380    -37     54   1.19   0.43    blue

178    -71     -5   -293    237   -113   -100   0.80   0.21     red
-1175     18    -81   -143   -239    -34     54   1.12   0.49    blue

 7    -24     51   -165    112     77   -100   0.84   0.13     red
-150   -105     63    131   -183   -492    -56   1.01   0.34    blue

894    -66     44     56    -52     77   -100   0.79   0.37     red
213    -48    153    -54    492   -326     -1   1.03   0.33    blue
tried to do Non-siderea tracking but it did not work - we sent suggestion to the PI how to adapt the OBs
 442    117      4   1323   -141   -294    -95   0.97   0.65     red
-441    -76     78    -39    -36    -67     49   1.31   0.20    blue

-332      9    -63    232     48     -2    -19   1.11   0.19     red
   -96    110    168   -128   -170    -52     50   1.43   0.15    blue

04:51 Slew to focus ob * lbcrangebal

04:54 Slew to WD_M35_N2158_LONG1
  • On the 1th images seeing is ~ 1.1-1.3
  • On the 2th images we are out of focus and out of collimation. Aborted!

05:13 dofpia,/X2
  • dimm seeing 1.4"
    4705   -164    -79    140    -23   -241   -100   0.70   1.86     red
    -394    146   -228     12   -293   -346    -45   0.88   0.32    blue
    
    -128   -205     30   -107    -27    122   -100   0.66   0.14     red
     -53     -2   -108   -201     13   -228    -20   0.94   0.19    blue
    
    601    -59     29    186    -55     18    -63   0.71   0.27     red
    -316     34   -106   -139    -24    -37     -8   1.04   0.17    blue
    
    158    101    -19    168    -85     51    -15   0.88   0.14     red
    612    -48     60    -41     10     17      9   1.11   0.26    blue
    

05:21 Slew to WD_M35_N2158_LONG1 again
  • Seeing meauserd on 1st images 1.44 blue and 1.28 red
  • 2nd images are elongated. Abort!

05:34 Slew to Donati: be27_focus_ok

05:36 dofpia,/X2
-2176     67     91   -209   -353     27     65   0.96   0.88     red
442     63   -108   -191   -164     23      7   1.36   0.23    blue

-1425    131    -87     15   -140    -62     23   0.76   0.57     red
504    -51    -73   -142    -23    -89    -15   1.13   0.23    blue

-696    152    -38    -34   -151    -59     16   0.64   0.30     red
140    217    -23     72     97     -3    -37   1.08   0.13    blue

1108     99    -29     15     30    -59     -1   0.70   0.45     red
 1739     16   -178     64    -31   -114    -59   1.07   0.70    blue

65     68    -21     42    -41     -6     11   0.65   0.09     red
234     -5    134    -86     26    -45    -33   1.06   0.14    blue

05:45 lbcrangebal

05:47 Slew to be27
  • images are elongated seeing on images 1.8 blue, 1.35 red
  • collimation doesn't work well.... we are waiting for termal stab of the mirrors
  • we will run ob again until mirrors reach termal stabilization

05:58 images seems to be no more elongated
  • seeing 1.69 blue 1.7 red

*06:04: Slew to be30_focus_ok

06:07 dofpia,/X2
  -610    -41    121     69   -369     17     26   0.80   0.31     red
 2891     -6    104    -66    -50   -132    -71   1.23   1.15    blue

 -807    -69     49    128   -113     66     20   0.63   0.34     red
  458    -28    108    -29    -52   -158    -59   1.05   0.22    blue

 930    -25     70   -150    -13    -38     18   0.80   0.38     red
1019    -37    -88    -64     43    -27    -25   1.25   0.41    blue

-1159     -7    -76      2    -80      1     20   0.66   0.47     red
 -867    -42     88     41     50     26    -26   1.08   0.35    blue

 794    -74     61    -29     18     36     25   1.10   0.32     red
506     63   -130     25    -48     43     12   1.29   0.22    blue

-855    -38    -10     27     22     35      6   0.75   0.35     red
-208     64    112     67    114    -39    -32   1.11   0.14    blue

 232    -31     24     13    -33     13      4   0.74   0.12     red
  447     25     94      9     17     49    -15   1.13   0.20    blue
  • lbcrangebal - during this 1sec exposure the red image is already out of focus!!!
  • dofpia again
 461    -29    -46    -64    -94   -149     -7   0.61   0.23     red
 759     51   -111    -52     12   -130    -38   0.99   0.32    blue

 918   -111    -13    -38     51     45      9   0.87   0.37     red
  1570    -44   -105    -56     21    -11    -16   1.17   0.63    blue

 -757     -1     52    -33    -89   -108     -6   0.60   0.32     red
 -495     98    130      9     33     -3    -23   1.02   0.22    blue

 -101     26    -54     23    -70    -27     -2   0.65   0.10     red
 -388    -65    -12    -49     23    -56    -13   1.01   0.18    blue

 -113    133     58     21    -15     19      1   0.63   0.10     red
 278    -23    -69    -54    -64    -56    -13   1.05   0.15    blue
 

06:30 be30
  • both chanels are elongated!

06:46 coll74 was too low...
  • dofpia
 1001    -69     33   -558    -71   -211    -38   0.66   0.49     red
-635    -10   -159   -256    272      3    -31   1.08   0.31    blue

 -189     80     34     94   -155     16    -23   0.70   0.14     red
 -49    -35   -111     21      7    -76    -19   1.04   0.10    blue

06:55 ngc2158_focus_ok
  • dofpia
 360    -84    -57    -39    263    -46     13   0.73   0.21     red
-178    -34    123    160    164     55     24   0.96   0.16    blue

 -391    109     28   -173    -97    -44      5   0.58   0.20     red
 -424     82    183    -25   -138     23     25   0.87   0.20    blue

 -56    -69    -51    180    -66    -88      6   0.58   0.13     red
 -206      6    157   -166     -3   -138     -8   0.91   0.17    blue
  • dimm seeing 1.4''
  • lbcrangebal

07:04 ngc2158
  • red image is way tooo elongated (0.27ellip) but the fwhm is <1.0''
  • the blue image is nice and round - fwhm <1.0''
  • this problem with the red channel has been going on for the last 2hrs!!!
  • red and blue seeing ~ 1" - but red images elongated!

07:21 gilli_focus1.ob
  • the seeing is good enough to try gilli now - if we still get problems with the collimation on the red side, there is nothing we can do...
  • dofpia,/X2
-34    144     73    332    -21   -205      7   0.55   0.22     red
-2528     -8    239   -502   -586    -56     50   0.94   1.06    blue

  -553    166      7   -451   -510   -127     35   0.59   0.40     red
 1031   -107    246    478    -52    -78     24   0.94   0.48    blue

  -95    -16    -54    466    339    -45     48   0.67   0.27     red
 482    143    201     90    -30   -114     30   0.83   0.23    blue

  582   -105     65   -252   -423   -152     19   0.80   0.34     red
-263   -153      6   -143     -7    -36     31   0.92   0.15    blue
  • lbcrangebal

07:33 gilli_ri1_1
  • seeing ~1", blue slightly elongated E=0.04
  • seeing is increasing in the blue channel 1.5", stable in the red 1"
  • seeing 1.46" blue , 1.1" red
  • seeing is go 1.4" bluw and 1.5" red

08:12 gilli_rz1-1.ob
  • dofpia,/X2
  • some problem occurs durig dofpia execution. Abort

08:21 dofpia again,/X2
4151    -34    167    -99   -293     68     32   0.83   1.64     red
-4947   -226   -157   -144    378    -56     18   0.77   1.96    blue

-14     52   -142   -308     92     17     52   0.71   0.17     red
-119      8    -99    -35    249     51     62   0.72   0.15    blue

-248     30    -84    157    186    -29     34   0.86   0.17     red
544    -98     93     -3     37    -50     34   1.16   0.23    blue

08:28 gilli_rz1-1.ob
  • Script aborted. Images are out of focus, bad collimation

08:39 dofpia again,/X2
570    135    -12   -413    -19    -62     23   0.83   0.31     red
354    -63   -121    -58     97   -132     25   0.82   0.19    blue

-792    -68    129    145    -87     52     31   0.84   0.34     red
533     21    -82   -182   -155   -202     25   0.90   0.28    blue

511     47     16    117   -136    -19      3   0.85   0.23     red
766     24   -128   -101   -170   -198     21   0.92   0.35    blue

168   -147    142     70    609    -60      4   0.93   0.30     red
237   -117   -180     98   -121   -107     20   0.96   0.17    blue

881    122    -53   -132    -60    -49      7   0.97   0.37     red
57    179     89   -256   -360     -9     28   1.33   0.22    blue

343    327   -229    254     48    -82      0   0.93   0.23     red
-781    -30    115    110    348     64     20   1.03   0.36    blue

08:51 dofpia aborted because of humidity 86.5%
  • We are reaching the dewpoint - close for humidity

-- KonstantinaBoutsia - 19 Feb 2012
Topic revision: r7 - 30 Mar 2012, RobertoSpeziali
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback